OpenStudy (anonymous):

Find cot θ if csc θ = sqrt 5/2 and tan θ > 0.

OpenStudy (anonymous):

\[\huge \cot^2 \theta = \csc^2 \theta -1 =(\frac{\sqrt 5}{2})^2 -1\] \[\huge =(\frac{5}{4}) -1= \frac {5-4}{4} =\frac{1}{4}\] \[\huge \rightarrow \cot \theta = \sqrt \frac{1}{4}= \frac{1}{2}\] @gamewarden23

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

@gamewarden23 medal please..